OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is 5x - 3 and whose width is x + 7? a. 5x2 - 21 b. 5x2 + 9x - 21 c. 5x2 + 32x - 21 d. 5x2 - 32x - 21

OpenStudy (jlg030597):

Im saying C

OpenStudy (jlg030597):

thats my guess cause im having a hard time concepting this stuff

hero (hero):

\[A_{\text{rect}} = ab\] Let a = 5x - 3 and b = x + 7 Then \[A_{\text{rect}} = (5x - 3)(x + 7)\] Multiply

OpenStudy (anonymous):

(5x-3)(x+7) 5x^2+35x-3x-21 5x^2+32x-21

hero (hero):

@jlg030597, this is not yahoo answers and you shouldn't help users with "guessing"

OpenStudy (jlg030597):

Ok sorry

hero (hero):

If any of what I say surprises you, I advise that you review the code of conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

5x^2+32x−21

OpenStudy (anonymous):

so c

OpenStudy (anonymous):